HUBBLE SPACE TELESCOPE DAILY REPORT #2311 PERIOD COVERED: 0000Z (UTC) 02/17/99 - 0000Z (UTC) 02/18/99 Daily Status Report as of 049/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.2 Completed WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 6588 (High Resolution Imaging of Ionized Gas in the Disk-Halo Interface of Spiral Galaxies) The WF/PC-2 was used to make high resolution imaging of the halo disk in spiral galaxy NGC-891 in order to study its ionized gas. Observations of the interstellar medium at the disk-halo interface of spiral galaxies provide us with very significant information on the physical state and structure of the interstellar medium in a global sense. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed STIS/MA1 7950 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to make a dark calibration image. This is to establish if the detectors are performing nominally, that no internal movement of the MCP has taken place, and establish dark level in orbit for a given operating temperature. These darks will be used in later image reduction and also trended over instrument life time as indicators of detector degradation or damage. The observations were completed as planned, and no problems were reported. 1.6 Completed Two Sets of STIS/CCD 8091 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude, non-scripted, parallel observations. The observations were completed as planned, and no anomalies were noted. 1.7 Completed STIS/CCD 7352 (The Nuclear Kinematics of Kinematically Distinct Cores) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the nuclear region of the kinematically distinct core of galaxy NGC-4365. A wave length calibration image was also taken. The observations were completed as planned, and no anomalies were reported. 1.8 Completed WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 1.9 Completed WF/PC-2 7377 (Dwarf Elliptical Galaxy Snapshot Survey II) The WF/PC-2 was used to make observations of the dwarf elliptical galaxy VCC963 as part of Snapshot Survey II. This survey will provide important information on the globular cluster systems of such galaxies and the properties of their cores and nuclei. The observations were completed as planned, and no anomalies were reported. 1.10 Completed WF/PC-2 7282 (A Continuing Snapshot Survey for Compact Stellar Groups and Clusters around Wolf-Rayet Stars) The WF/PC-2 was used to make observations of the Wolf-Rayet star WR112 as part of a continuing snapshot survey of compact stellar groups and clusters. The newly imaged clusters allow us to accurately determine the distances, luminosities, and ages of the WR members. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 15 Successful : 15 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared once. Per HSTAR 7005, the recharge ratios for batteries 5 and 6 went out-of-limits high at 048/155315Z. The +E SPA was disconnected at 048/1830Z as directed by an operations request. At approximately 048/2049Z, the recharge ratios went back in bounds. At 048/211415Z, the NICMOS FPA temperature sensor on detector 2 disabled when the temperature monitor had a value in excess of 300 degrees K. Per an operations note, an existing operations request was executed at 049/0128Z which re-enabled the FPA temperature sensor. A TTR was written to document data hits and missing minor frames at 049/072808Z and 049/084155Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH